If y varies directly with x, and y = 500 when x = 200, find the constant of variation k.
Answer by ankor@dixie-net.com(22740)

"If y varies directly with x,"
y = x*k
:
"and y = 500 when x = 200, find the constant of variation k."
Substitute for y and x:
500 = 200k
k = 500/200
k = 2.5
:
Check:
y = 2.5(200)
y = 500